a. Work through the problem below and show working for all variables (1-10) in addition to
totals of columns and rows in the matrix. (10 marks)
C = 300 + 0.8YD (Private consumption)
YD = Y + TR – T (Private Disposable income)
T = 80 + 0.1Y (Total taxes)
G = 300, TR = 100 (Gov. expenditure and transfer payments respectively)
I = 450 (Private investment)
X = 200 (Total exports)
IM = 120 + 0.2Y (Total imports)
Y = C + I + G + X – M (Goods Market Equilibrium condition)
Policy variables: Fiscal policy (G, t and TR), Monetary Policy: (N/A)
1
Expert's answer
2019-05-13T09:35:54-0400

Y = C + I + G + X – M = 300 + 0.8×(Y + 100 - 80 - 0.1Y) + 450 + 300 + 200 - 120 - 0.2Y,

Y - 0.72Y + 0.2Y = 300 + 16 + 450 + 300 + 200 - 120,

0.48Y = 1146,

Y = 2387.5.